# HAMLOCK: HArdware-Model LOgically Combined attacK

### v1

@@ 7 Conclusion
Appendix A Outline
As part of the Appendix for "HAMLOCK: HArdware-Model LOgically Combined attacK", we provide additional details and experiment results as organized below:
  • Appendix˜B details the model-level white-box and black-box defenses evaluated in the main paper.
  • Appendix˜C presents further experimental results, including activation value distributions and example backdoor triggers across benchmark datasets.
  • Appendix˜D offers additional background on hardware Trojan attacks, including implementation strategies on FPGA and ASIC platforms.
Appendix B Details on Model Level Defenses
In this section,we provide detailed descriptions of the model-level defenses evaluated in our study. Section˜B.1 discusses defenese that detect backdoored models, Section˜B.2 outlines backdoor sample detection techniques, Section˜B.3 covers black-box backdoor sample detection techniques, and Section˜B.4 describes model hardening approaches aimed at mitigating the backdoor effects.
B.1 Detecting Backdoored Models
These defenses check whether a given model is backdoored at the model level.
Neural Cleanse (NC): NC attempts to reverse-engineer potential backdoor triggers for all possible target classes by identifying minimal perturbations that can induce misclassification. It then computes an anomaly index to detect whether any class requires significantly smaller triggers than others. If such a case exists, the model is flagged as backdoored. NC is particularly effective against small, patch-based triggers but becomes ineffective when the attacker uses large or dispersed trigger patterns, as its trigger-recovery procedure is biased toward minimal triggers [23].
MNTD: The MNTD defense [65] trains a meta-classifier to distinguish between clean and malicious models. To do this, it first creates a large dataset of “shadow" models, consisting of both benign models and models backdoored with a variety of known attack methods. The meta-classifier is then trained on this dataset to learn the generalizable statistical patterns in neuron weights that are characteristic of a backdoor. When inspecting a new candidate model, MNTD uses this learned knowledge to determine if it is clean or has been compromised.
B.2 Backdoor Sample Detection
These defenses check whether a given input sample contains a trigger. We evaluate two recent state-of-the-art techniques:
TED: TED treats a neural network as a dynamic system in which inputs evolve through intermediate feature representations toward final predictions. It analyzes the trajectory of class-specific nearest neighbors across layers. Clean samples follow consistent paths, while backdoor samples deviate. The variance of the class-rank trajectory serves as a detection signal. Higher variance indicates anomalous behavior. Detection is performed using unsupervised outlier detection frameworks such as PyOD [^1] https://github.com/yzhao062/pyod.
IBD-PSC: This defense exploits the observation that backdoored samples tend to produce more consistent predictions under model parameter amplification. For a given input, the model parameters are scaled with various amplification factors, and the variance in confidence scores is analyzed. Backdoor samples exhibit lower variance and higher average confidence compared to clean inputs. IBD-PSC requires batch normalization layer for the victim model and hence, we did not include the results on MNIST in the main paper, as LeNet model does not have batch normalization.
B.3 Black-box Backdoor Detection Methods
Besides white-box backdoor sample detection methods, there also exists black-box detection methods that only require the input output information. We describe the classical STRIP [16] as well as the state-of-the-art detection BBCal [26] below.
STRIP: STRIP overlays a given input with a set of randomly selected clean samples and observes the consistency of model predictions. If predictions remain unchanged under perturbation (i.e., low output entropy), the sample is likely to contain a backdoor trigger. High entropy indicates normal (clean) behavior.
BBcal: BBcal leverages causal analysis and distinguishes between clean and backdoor samples by measuring the break point where predictions change. Backdoor samples either have very low or high break point, which is then leveraged to filter out the backdoored samples.
B.4 Backdoor Mitigation
These defenses mitigate the effects of a backdoor by lightly retraining the model rather than detecting backdoors only.
Fine-tuning and fine-pruning: Fine-tuning [49] remove the backdoor effectiveness by tuning the backdoored model on a set of clean training data. Fine-pruning [36] removes low-activation neurons by setting their weights to 0. These methods focus on suppressing backdoor behavior without damaging overall model accuracy.
BEAGLE: BEAGLE builds on fine-tuning and performs forensic analysis to mitigate a wider variety of backdoored models. It assumes access to a small number of backdoor samples and a larger clean dataset. In our setup, we evaluate BEAGLE in the worst-case scenario for the attacker, where the defense uses backdoor samples directly extracted from the same model being inspected, while in the original paper, such samples are often obtained from other backdoored models.
Figure 4: The activation value distributions of clean and backdoor samples
Figure 5: Backdoor triggers on the three benchmark datasets. The triggers are the squared regions on each corner of different images. The top row shows trigger samples used for the trigger optimization attack, while the bottom row displays trigger samples generated by the weight optimization attack.
@@ Appendix C Additional Results
In Figure˜4, we show the distribution of activation values in clean and backdoor samples for the CIFAR10 ResNet model. We can clearly see a nonzero threshold that separates the clean and backdoor activations, and such a threshold can be obtained by slightly tuning the scaling factor $s$.
@@
Figure 6: An example of a System-on-Chip (SoC) containing an AI accelerator with three types of hardware Trojans: weight Trojan, trigger Trojan, and payload Trojan. The weight Trojan is integrated with the memory subsystem and perturbs selected model weights to instantiate trigger neurons. The trigger Trojan monitors the output activity of these trigger neurons and logs temporal patterns associated with input triggers. The payload Trojan remains dormant until activated by the trigger Trojan, at which point it stealthily perturbs selected output neurons to induce misclassification.
Appendix D Background on Hardware Trojan Attacks
HAMLOCK works by monitoring the distinct activation on the trigger neuron embedded inside the model using trigger hardware Trojan and then enabling the payload hardware Trojan accordingly. The different types of hardware Trojans are illustrated in Figure 6. Below, we describe how the trigger Trojan and payload Trojan are implemented on different hardwares.
Trigger Trojan Logic. This hardware Trojan unit monitors the abnormal activation of a specific neuron—designated at the model level as the trigger neuron. On FPGAs, this monitoring is implemented by intercepting memory accesses to the address corresponding to the trigger neuron’s output. On ASICs, the neuron’s datapath is directly tapped. To optionally introduce a temporal constraint—e.g., requiring the trigger pattern to occur only after certain number of inferences—compact comparator and counter logic are included to track the neuron’s activation over time. If the trigger neuron exhibits its characteristic backdoor activation for a predefined number of $N$ occurrences, the Trojan asserts a payload-enable signal. This mechanism ensures that the Trojan remains dormant until high confidence is achieved that a backdoor input is present.
Payload Trojan Logic. Once the payload-enable signal is enabled, the payload Trojan executes its function by perturbing the final logit values corresponding to a specific target class in the interest of the attacker. This manipulation is not performed at the model level but rather injected directly into the hardware datapath. For FPGAs, this may involve injecting a bias into the MAC computation path or inserting a fixed perturbation into the fetched value of the logit neurons. For ASICs, where datapath access is more direct, the payload Trojan slightly increases the activation of a payload neuron or directly alters the class scores before softmax. This ensures that, even though the model has predicted correctly in the logic level, the final output—e.g., what is sent to the host or printed by the accelerator—is corrupted. The misclassification is therefore the result of a hardware-induced misrepresentation of the logit space, not a flaw in the model itself.
Weight Modification Trojan. Beyond trigger and payload Trojans, our co-attack framework can also incorporate a weight modification Trojan to directly alter model parameters at the hardware level without modifying the golden model parameters. This type of Trojan is particularly useful in scenarios where the attacker has hardware-level access but cannot modify the model weights directly in software—such as in off-the-shelf or third-party deployed systems (see Section˜2.2). The weight Trojan enables subtle tampering of the model’s internal structure to strengthen or activate backdoor behavior, while remaining invisible to traditional software-based validation or retraining.
At a high level, the Trojan selectively perturbs specific weights associated with the target trigger neuron, either by modifying the fetched value at runtime or by rewriting the memory during initial deployment. These perturbations are designed to enforce or enhance the abnormal activation behavior required by the subsequent trigger Trojan logic. This can be done through adversarial logic embedded directly in the memory access path. By embedding the weight modification at the hardware level, the attacker ensures that the model’s parameters appear clean under inspection but behave maliciously when executed in real devices.
With the description of the role of trigger Trojan and payload Trojan, we next describe in detail about the actual implementation of these two types of trojans on FPGAs and ASICs.
Details on FPGA Platforms. On FPGA platforms, the designs are typically resource constrained and relies heavily on reutilizing existing logic elements. Hence, it is highly likely that the model parameters would be stored in a centralized memory subsystem, where all model weights and activations are stored and accessed through a shared address space. Thus different variations of the Trojans can be attached to different components of the said AI Accelerator. The trigger Trojan can be realized by inserting lightweight logic into the output channels of logic elements that are responsible for calculating the neuron activations. It monitors the neuron IDs of the ongoing operations and if it matches the trigger neurons, it compares the computed activation value to a predefined threshold. To introduce temporal stealth, a compact counter can be used to track repeated activations across inference runs. When the trigger condition is satisfied, a control signal is raised to activate the payload Trojan logic.
The payload Trojan can be realized by inserting the Trojan to the input channels of logic elements that are responsible for calculating the neuron activations. Once enabled, it perturbs logit values when payload Neuron IDs are encountered. These perturbations are injected by altering data in-flight to the MAC units or applying additive bias before the softmax computation. Crucially, the manipulation is transient and does not alter stored model parameters, allowing the model to pass offline evaluations unmodified.
The optional weight modification Trojan can be attached to the memory subsystem itself. It selectively modifies weight values either during fetch (transient manipulation) or during initial access (persistent modification). When specific weight addresses are accessed, typically those associated with the trigger neuron, the Trojan injects perturbed values on the data bus while keeping the underlying memory content untouched. This allows the attacker to fine-tune the model’s behavior at deployment time without retraining or modifying the model file.
Details on ASIC Platforms. On ASIC platforms, the design may be less resource constrained and hence there can have greater hardware unpacking. The trigger Trojan can be implemented by snooping the output channel of the trigger neuron and monitoring its output values in real time. The Trojan can be a simple comparator and counter logic can be directly embedded into the neuron’s computation block, allowing for ultra-compact and efficient detection of backdoor activation patterns.
The payload Trojan can be similarly embedded near the final classification layer. When a Trojan payload enable signal is asserted by the trigger logic, it perturbs either the activation of a high-sensitivity neuron or directly modifies the logit output before classification. Given the flexibility in ASIC design, the perturbation signal can be routed through existing or non-critical channels to avoid suspicious signal fan-outs or layout anomalies, enhancing stealth.
The weight modification Trojan in ASICs can be inserted close to the memory element that stores all of the weights and biases. It modifies the selected weights whenever they are accesssed. Perturbations may include bit-flips, scalar shifts, or noise injections. These modifications allow the attacker to enforce or amplify the backdoor effect from the model-level attack. Because ASIC designs provide fine-grained placement and routing control, the Trojan logic can remain well hidden with existing logic elements or unused silicon space.
In both platforms, the three types of Trojans can operate in coordination: the weight Trojan optionally sets up the backdoor behavior; the trigger Trojan monitors runtime conditions; and the payload Trojan corrupts the final output. Their physical separation, minimal footprint, and context-sensitive activation make this co-attack stealthy, persistent, and difficult to detect through conventional software-based verification.

### v2

@@ 7 Conclusion
8 Acknowledgement
We thank the anonymous reviewers and the shepherd for their insightful comments and constructive suggestions, which significantly improved this work. Fnu Suya acknowledges support from the AI TechX Seed Funding Program at the University of Tennessee, Knoxville. Swarup Bhunia is supported by National Science Foundation under Grant No 2350365.
Appendix A Ethical Considerations
We introduce a novel hardware–software combined backdoor attack with the goal of exposing critical vulnerabilities in the deployment of AI models on hardware accelerators. We consider the ethical implications of our research as follows.
Stakeholders and Potential Impact. The stakeholders potentially affected by this work include: 1) Hardware vendors: designers and integrators of AI accelerators, IP vendors, and EDA tool providers; 2) Model distribution platforms: repositories and marketplaces that host pre-trained or third-party models; 3) Edge AI/IoT system users: organizations and end users deploying ML workloads on accelerators; and 4) Security research communities: researchers, practitioners, and policymakers focused on AI security and supply-chain integrity. This work may have varying impact on these groups. For hardware vendors, this work highlights a risk scenario involving compromised design flows or third-party components. For model distribution platforms, our findings illustrate that software-only validation may be insufficient when models interact with untrusted hardware, motivating stronger provenance checks and deployment-time safeguards. For system users, misuse of such attacks could lead to degraded or incorrect system behavior in safety- or performance-critical settings. For the research community, this research expands the threat model for ML supply chains and creates urgency for cross‑layer defenses that combine hardware security, software verification, and ML robustness.
Mitigation of Negative Impacts. To minimize potential harm, all experiments were conducted under controlled conditions using publicly available, non-sensitive datasets. We emphasize that this work is intended for defensive research, aiming to responsibly identify an unexplored hardware–model attack surface, and we explicitly discourage any malicious use. HAMLOCK’s co-design represents a novel class of vulnerability for which no existing off-the-shelf defenses currently exist. Effective mitigation will likely require new research into hardware–software co-verification. For example, future EDA tools could cross-reference a compiled model’s datapath with the hardware layout to detect potential anomalies. Long-term, actionable mitigations include promoting hardware–software co-verification, strengthening supply-chain vetting, and increasing transparency in EDA toolchains. Because this vulnerability arises from the custom ASIC life cycle, we are coordinating with our institution to share these findings with the broader hardware security community. Upon publication, we also plan to engage with major EDA tool vendors (e.g., Synopsys, Cadence) and IP providers, as their tools/designs are well-suited for exploring the co-verification defenses.
Decision Rationale. We conducted this research to study the underexplored hardware–software interface, motivating proactive defenses across AI and hardware security communities.
Appendix B Open Science
We have made all code publicly available at https://zenodo.org/records/18022955. All datasets used are publicly accessible through their official sources.
@@ Appendix C Additional Results
In Figure˜4, we show the distribution of activation values in clean and backdoor samples for the CIFAR10 ResNet model. We can clearly see a nonzero threshold that separates the clean and backdoor activations, and such a threshold can be obtained by slightly tuning the scaling factor $s$ in Eq. (3).
@@
Fine-Tuning with NAD and DT. Beyond defenses covered in Section˜5.2, we also evaluated Neural Attention Distillation (NAD) [36] and Democratic Training (DT) [55]. NAD is a defense previously shown to underperform compared to our evaluated baseline BEAGLE [9]. DT, while originally designed for universal adversarial perturbations, was assessed for its potential applicability against backdoor attacks. The results, presented in Table˜9, demonstrate that both defenses have almost no impact on the attack effectiveness. Specifically, two variants of the 1-N attack retained 100% effectiveness against both NAD and DT. The more generalized 3-N attack saw only a slight drop in effectiveness (around 1%) for GTSRB and CIFAR10. On ImageNet, DT performed slightly better, reducing the ASR from 97% before the defense to 90% afterward, yet this came at the cost of a 6% drop in clean accuracy. More importantly, an ASR of 90% still indicates effectiveness of the defense in practice.
Table 9: Effectiveness of HAMLOCK against Neural Attention Distillation and Democratic Training on different datasets. The model architecture is ResNet-18. “NDA” denotes Neural Attention Distillation, “DT” denotes Democratic Training.
Datasets Attack ASR (%) Clean Acc (%) ASR (%)
w/o Defense NDA DT NDA DT
GTSRB 1N-trigger 100.0 93.5 89.4 100.0 100.0
1N-weight 100.0 93.5 89.4 100.0 100.0
3N-weight 97.0 96.1 92.2 95.7 95.2
CIFAR10 1N-trigger 100.0 91.3 89.9 100.0 100.0
1N-weight 100.0 91.4 90.0 100.0 100.0
3N-weight 96.0 90.8 90.7 95.4 94.8
Imagenet 1N-trigger 100.0 65.0 58.6 100.0 100.0
1N-weight 100.0 65.2 59.2 100.0 100.0
3N-weight 97.0 65.0 59.5 95.4 90.1
Appendix D Algorithm
Algorithm 1 outlines the HAMLOCK multi‑neuron weight optimization process described in Section˜3.2.2.
Algorithm 1 Multi-Neuron Weight Optimization
1:Input: Pre-trained Model $f_{\theta}$, Trigger $\delta$, Scaling Factor $s$, Number of Layers $r$, Neurons per Layer $k$, Calibration Set $\mathcal{D}_{cal}$ ($|D_{cal}|$ = $M$), Accuracy Drop Tolerance $\tau$
2:Output: Backdoored Model $f_{\theta^{\prime}}$
[…]
4:Randomly select $r$ layers from $f_{\theta}$. $\triangleright$ Layer Section
5:for each selected layer $l$ do
6:  $\mathcal{C}\leftarrow\emptyset$
7:  for each neuron $j$ in layer $l$ do
8:   Temporarily set $a_{j}(x)\leftarrow 0$ for all $x\in\mathcal{D}_{cal}$
9:   Measure accuracy drop $\Delta_{\text{acc}}$
10:   if $\Delta_{\text{acc}}<\tau$ then
11:     $\mu_{clean}\leftarrow\frac{1}{M}\sum a_{j}(x)$
12:     $\mu_{bd}\leftarrow\frac{1}{M}\sum a_{j}(x^{\prime})$
13:     $S_{j}\leftarrow|\mu_{bd}-\mu_{clean}|$
14:     $\mathcal{C}\leftarrow\mathcal{C}\cup\{(j,S_{j})\}$
15:   end if
16:  end for
17:  $\mathcal{T}\leftarrow\operatorname{TopK}(\mathcal{C},k)$ $\triangleright$ Candidate Selection
18:  for each target neuron $j\in\mathcal{T}$ do
19:   $w_{j}$: weights connecting layer l1l-1 to neuron $j$.
20:   for each connection $i$ from layer l1l-1 do
21:     $w^{\prime}_{ji}\leftarrow s\cdot|w_{ji}|\cdot\text{sgn}(\bar{a_{i}}(x^{\prime})-\bar{a_{i}}(x))$
22:   end for
23:   Update neuron $j$ weights in $f_{\theta}$ with $W^{\prime}_{j}$.
24:  end for
25:end for
26:return $f_{\theta}$